# Theory of Irreversibility in Quantum Many-Body Systems

### v2

@@ III.4.2 Decay of the OTOC in the DRPM
for $x>y$. Similarly to the case of PBCs, we already know that the connected part decays with the rate $2\varepsilon$ for any value of $y$ at late times while $L$ is fixed. This turns out to be not the case for the norm $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=\sum_{s,s^{\prime}}[\widehat{S}_{\mathrm{d}}^{y-1}]_{s0}\omega_{s^{\prime}}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s^{\prime}}$, which is sensitive to the initial location of the operator $\mathcal{O}(y,0)$ for OBCs. It can be readily seen that when $y$ is sufficiently close to the boundary $y=1$ (resp. $y=L)$ such that $\lim_{L\to\infty}y-1<\infty$ (resp. $\lim_{L\to\infty}L-y<\infty$), we have $\lim_{\omega\to 1}\lim_{L\to\infty}[\widehat{S}_{\mathrm{d}}^{y-1}]_{s0}=1$ (resp. $\lim_{\omega\to 1}\lim_{L\to\infty}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s}\omega_{s}=1$), yielding
$\lim_{\omega\to 1}\lim_{L\to\infty}\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=\lim_{\omega\to 1}\lim_{L\to\infty}\sum_{s}\omega_{s}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s}$ (66)
@@ IV Suppression of anomalous relaxation
Having studied in detail the relaxation dynamics of the simplest version of the DRPM, in this section, we inquire what the conditions for observing anomalous relaxation in quantum many-body systems are. To this end, we study minimal extensions that alter some of the properties in the DRPM, namely, extensive (bulk) dissipation, absence of conservation laws, ergodicity, and local interactions. We shall show that in the absence of these properties, anomalous relaxation is suppressed, i.e., certain operators do not relax, and the gap extracted from the DFF is not positive finite when the dissipationless limit is taken after the thermodynamic limit.
@@ IV.1 Subextensive dissipation
where $m=1,\dots,M$ enumerates the clusters of sites on which quantum channels do not act, and ¯$\bar{\ell}_{m}$ is the length of the cluster labelled by $m$. This indicates that the gap is given by
@@ IV.2 Nonlocal interactions
Next, we consider a chaotic quantum circuit with fully nonlocal interactions. We define the random matrix Floquet (RMF) system as a Floquet system whose time-evolution operator $W$ is a $q^{L}\times q^{L}$ matrix drawn from the circular unitary ensemble, and let us call the RMF system with local dissipation the dissipative RMF (DRMF) system. The DFF of the DRMF system is then defined in the usual way,
@@ IV.3.1 RPM with a conserved $U(1)$ charge
$\overline{K_{N}(t)}=\overline{({\rm Tr}_{N}W^{t}({\rm Tr}_{N}(W^{\dagger})^{\dagger})}=2^{-L}D_{N}\sum_{\underline{\alpha}}\overline{C^{[N]}_{\underline{\alpha}\underline{\alpha}}(t)},$ (80)
@@ IV.3.2 DRPM with a conserved $U(1)$ charge
Figure 7: Numerical spectral gap of qubit brickwork RFCs as a function of dissipation strength $\gamma$ for different system sizes. The Haar-random RFC (top left) and Heisenberg RFC at strong coupling (top right) are quantum chaotic and show the finite-size signatures of anomalous relaxation. The RFC with U(1) conservation (bottom left) and the Heisenberg RFC in the localized phase (bottom right) do not show anomalous relaxation, as the gap closes in the thermodynamic limit at small $\gamma$. The spectral gap was computed by a power iteration method [17] for the Haar and Heinsenbergs RFCs and by ED for the U(1) RFC. We have checked that for system sizes amenable to ED, the gaps obtained by power iteration and ED coincide for the Haar RFC until $\gamma\approx 0.25$ (above which there is a small systematic deviation of about $3\%$) and for the Heisenberg RFC with $J=0.01\pi$ at all $\gamma$. For the U(1) circuit, there was a large systematic deviation, hence we resorted only to the ED-computed gap.
@@ IV.4 Numerical results for $q=2$
where $u_{1,2,3,4}$ are independent $q\times q$ Haar-random unitaries, $J$ is the nearest-neighbor interaction strength, and $\mathbb{P}_{q}$ is the two-site Swap gate. At $q=2$, Eq. (93) is equivalent to a Trotterized Heisenberg interaction in random fields, and by abuse of nomenclature, refer to it as a Heisenberg gate for arbitrary $q$. For small enough $J$, neighboring sites are weakly coupled and the system localizes, while for larger $J$ the interactions dominate and we enter an ergodic phase (for $q=2$, the transition occurs at $J_{c}\approx 0.07\pi$ [86]). We considered the Heisenberg RFC in the ergodic phase for $J=0.3\pi$ and in the localized phase for $J=0.01\pi$.
To compute the spectral gap at finite dissipation strength, we employed exact diagonalization (ED) and a power iteration method described in Ref. [17]. In Fig. 7 we show the numerical results for the spectral gap computed for the four different qubit brickwork RFCs and different system sizes as a function of dissipation strength. For the ergodic circuits (the Haar RFC and the Heisenberg RFC at large $J$), we see the finite-size signatures of anomalous relaxation (i.e., a nonvanishing leading RP resonance): although the gap must close at any finite system size in the $\gamma\to 0$ limit, the slope at small $\gamma$ is increasing with $L$. At larger $\gamma$, curves for different $L$ collapse to a single one, which persists until $\gamma$ is small enough to probe the finite spacings between individual eigenvalues. Accordingly, as $L$ increases, the single collapse curve extends until smaller values of $\gamma$, and, in the thermodynamic limit, it extends to $\gamma=0$, consistent with a diverging slope at the origin in this limit.
@@ Appendix A Partial dissipative form factor
While a closed expression of Eq. (96) is still not feasible to have, its late-time behavior can be inferred by the domain-wall expansion as before, yielding
@@ Appendix B Systems with nonlocal interactions
As in the DRPM, upon Haar averaging, only cyclic pairings of indices contribute at large $q$, which we parameterize as $(a(\tau),b(\tau))=(a^{*}(\tau+s),b^{*}(\tau+s))$ for $s=0,\dots,t-1$. As a result, for example at $t=2$, we obtain the large-$q$ DFF

### v3

@@ III.4.2 Decay of the OTOC in the DRPM
for $x>y$. Similarly to the case of PBCs, we already know that the connected part decays with the rate $2\varepsilon$ for any value of $y$ at late times while $L$ is fixed. This turns out to be not the case for the norm $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=\sum_{s,s^{\prime}}[\widehat{S}_{\mathrm{d}}^{y-1}]_{s0}\omega_{s^{\prime}}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s^{\prime}}$, which is sensitive to the initial location of the operator $\mathcal{O}(y,0)$ for OBCs. It can be readily seen that when $y$ is sufficiently close to the boundary $y=1$ (resp. $y=L)$ such that $\lim_{L\to\infty}y-1<\infty$ (resp. $\lim_{L\to\infty}L-y<\infty$), we have $\lim_{\omega\to 1}\lim_{L\to\infty}[\widehat{S}_{\mathrm{d}}^{y-1}]_{s0}=1$ (resp. $\lim_{\omega\to 1}\lim_{L\to\infty}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s}\omega_{s}=1$), yielding
$\lim_{\omega\to 1}\lim_{L\to\infty}\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=\lim_{\omega\to 1}\lim_{L\to\infty}\sum_{s}\omega_{s}[\widehat{S}_{\mathrm{d}}^{L-y}]_{0s}$ (66)
@@ IV Suppression of anomalous relaxation
Having studied in detail the relaxation dynamics of the simplest version of the DRPM, in this section, we inquire what the conditions are for observing anomalous relaxation in quantum many-body systems. To this end, we study minimal extensions that alter some of the properties in the DRPM, namely, extensive (bulk) dissipation, absence of conservation laws, ergodicity, and local interactions. We shall show that in the absence of these properties, anomalous relaxation is suppressed, i.e., certain operators do not relax, and the gap extracted from the DFF is not positive finite when the dissipationless limit is taken after the thermodynamic limit.
@@ IV.1 Subextensive dissipation
where $m=1,\dots,M$ enumerates the clusters of sites on which quantum channels do not act, and ¯$\bar{\ell}_{m}$ is the length of the cluster labeled by $m$. This indicates that the gap is given by
@@ IV.2 Nonlocal interactions
Next, we consider a chaotic quantum circuit with fully nonlocal interactions. We define the random matrix Floquet (RMF) system as a Floquet system whose time-evolution operator $W$ is a $q^{L}\times q^{L}$ matrix drawn from the circular unitary ensemble, and we call the RMF system with local dissipation the dissipative RMF (DRMF) system. The DFF of the DRMF system is then defined in the usual way,
@@ IV.3.1 RPM with a conserved $U(1)$ charge
$\overline{K_{N}(t)}=\overline{({\rm Tr}_{N}W^{t})({\rm Tr}_{N}(W^{\dagger})^{t})}=2^{-L}D_{N}\sum_{\underline{\alpha}}\overline{C^{[N]}_{\underline{\alpha}\underline{\alpha}}(t)},$ (80)
@@ IV.3.2 DRPM with a conserved $U(1)$ charge
Figure 7: Numerical spectral gap of qubit brickwork RFCs as a function of dissipation strength $\gamma$ for different system sizes. The Haar-random RFC (top left) and Heisenberg RFC at strong coupling (top right) are quantum chaotic and show the finite-size signatures of anomalous relaxation. The RFC with U(1) conservation (bottom left) and the Heisenberg RFC in the localized phase (bottom right) do not show anomalous relaxation, as the gap closes in the thermodynamic limit at small $\gamma$. The spectral gap was computed by a power iteration method [17] for the Haar and Heisenberg RFCs and by ED for the U(1) RFC. We have checked that for system sizes amenable to ED, the gaps obtained by power iteration and ED coincide for the Haar RFC until $\gamma\approx 0.25$ (above which there is a small systematic deviation of about $3\%$) and for the Heisenberg RFC with $J=0.01\pi$ at all $\gamma$. For the U(1) circuit, there was a large systematic deviation, hence we resorted only to the ED-computed gap.
@@ IV.4 Numerical results for $q=2$
where $u_{1,2,3,4}$ are independent $q\times q$ Haar-random unitaries, $J$ is the nearest-neighbor interaction strength, and $\mathbb{P}_{q}$ is the two-site Swap gate. At $q=2$, Eq. (93) is equivalent to a Trotterized Heisenberg interaction in random fields, and with an abuse of nomenclature, we refer to it as a Heisenberg gate for arbitrary $q$. For small enough $J$, neighboring sites are weakly coupled and the system localizes, while for larger $J$ the interactions dominate and we enter an ergodic phase (for $q=2$, the transition occurs at $J_{c}\approx 0.07\pi$ [86]). We considered the Heisenberg RFC in the ergodic phase for $J=0.3\pi$ and in the localized phase for $J=0.01\pi$.
To compute the spectral gap at finite dissipation strength, we employed exact diagonalization (ED) and a power iteration method described in Ref. [17]. In Fig. 7 we show the numerical results for the spectral gap computed for the four different qubit brickwork RFCs and different system sizes as a function of dissipation strength. For the ergodic circuits (the Haar RFC and the Heisenberg RFC at large $J$), we see the finite-size signatures of anomalous relaxation (i.e., a nonvanishing leading RP resonance): although the gap must close at any finite system size in the $\gamma\to 0$ limit, the slope at small $\gamma$ is increasing with $L$. At larger $\gamma$, curves for different $L$ collapse to a single one, which persists until $\gamma$ is small enough to probe the finite spacings between individual eigenvalues. Accordingly, as $L$ increases, the single collapsed curve extends until smaller values of $\gamma$, and, in the thermodynamic limit, it extends to $\gamma=0$, consistent with a diverging slope at the origin in this limit.
@@ Appendix A Partial dissipative form factor
While a closed expression of Eq. (96) is still not feasible, its late-time behavior can be inferred by the domain-wall expansion as before, yielding
@@ Appendix B Systems with nonlocal interactions
As in the DRPM, upon Haar averaging, only cyclic pairings of indices contribute at large $q$, which we parametrize as $(a(\tau),b(\tau))=(a^{*}(\tau+s),b^{*}(\tau+s))$ for $s=0,\dots,t-1$. As a result, for example at $t=2$, we obtain the large-$q$ DFF